# On estimates of trace-norm distance between quantum Gaussian states

### v3

@@ 1 Introduction and results
$\left\|\rho_{m_{1},\alpha}-\rho_{{}_{m_{2},\beta}}\right\|_{1}\leq\sqrt{16\min\left(a,b\right)\left\|m_{1}-m_{2}\right\|_{2}^{2}+\left\|\alpha-\beta\right\|_{2}^{2}},$ (1)
@@
$2\sqrt{8\min\left(a,b\right)\left\|m_{1}-m_{2}\right\|_{2}^{2}+\frac{1}{2}\left\|\alpha-\beta\right\|_{2}^{2}+\min\left\{P(a,b),P(b,a)\right\}\left\|\alpha-\beta\right\|_{1}},$
@@ A Gaussian state $\rho_{m,\alpha}$ is determined by its quantum characteristic function
$\mathrm{Tr}\sqrt{\rho_{\alpha}}\sqrt{\rho_{m,\beta}}=\frac{\left(\det\hat{\alpha}\det\hat{\beta}\right)^{1/4}}{\det\sigma(\hat{\alpha},\hat{\beta})^{1/2}}\exp\left(-\frac{1}{2}m^{t}\sigma(\hat{\alpha},\hat{\beta})^{-1}m\right),$ (29)
@@
$\exp\left[-\frac{1}{4}\mathrm{Tr\,}d(\alpha,\beta)-m^{t}\left(\hat{\alpha}+\hat{\beta}\right)^{-1}m\right]\leq\mathrm{Tr}\sqrt{\rho_{\alpha}}\sqrt{\rho_{m,\beta}}$ (30)
$\leq$ $\exp\left[-\frac{1}{4}\frac{\ln(1+\left\|d(\alpha,\beta)\right\|)}{\left\|d(\alpha,\beta)\right\|}\mathrm{Tr\,}d(\alpha,\beta)-m^{t}\left(\hat{\alpha}+\hat{\beta}\right)^{-1}m\right].$ (31)
@@
$1-\left(\mathrm{Tr}\sqrt{\rho_{\alpha}}\sqrt{\rho_{m,\beta}}\right)^{2}\leq 2m^{t}\left(\hat{\alpha}+\hat{\beta}\right)^{-1}m+\frac{1}{2}\mathrm{Tr\,}d(\alpha,\beta).$
@@
$\left\|\rho_{\alpha}-\rho_{{}_{m,\beta}}\right\|_{1}\leq 2\sqrt{2m^{t}\left(\hat{\alpha}+\hat{\beta}\right)^{-1}m+\frac{1}{8}\mathrm{Tr}\left(\hat{\alpha}-\hat{\beta}\right)\left(\hat{\beta}^{-1}-\hat{\alpha}^{-1}\right)}.$ (32)
@@ 3.1 Pure states
$\left\|\rho_{\alpha}-\rho_{{}_{m,\beta}}\right\|_{1}\leq 2\sqrt{m^{t}\left(\alpha+\beta\right)^{-1}m+\frac{1}{4}\mathrm{Tr}\left(\alpha-\beta\right)^{2}}.$
@@
$\left\|\rho_{\alpha}-\rho_{{}_{m,\beta}}\right\|_{1}\leq\sqrt{16\min\left(a,b\right)\left\|m\right\|_{2}^{2}+\left\|\alpha-\beta\right\|_{2}^{2}},$ (38)
@@ 3.2 Gauge-invariant states
$\frac{1}{8}\mathrm{Tr}\left(\hat{\alpha}-\hat{\beta}\right)\left(\hat{\beta}^{-1}-\hat{\alpha}^{-1}\right)\leq\frac{1}{2}\left\|\alpha-\beta\right\|_{2}^{2}+\frac{1}{2}\left\|\alpha\Upsilon_{\alpha}-\beta\Upsilon_{\beta}\right\|_{2}^{2}.$ (44)
@@
$=$ $P(a,b)\left\|\beta-\alpha\right\|_{1}$
By using (44) we obtain the estimate (3) for the trace-norm distance of two Gaussian states.
@@ 5 Appendix: the case of fermionic Gaussian states of CAR
and it is called covariance operator of the state $\rho\left(\cdot\right).$ Conversely, any such operator defines a quasifree state according to (49), (48) which is denoted $\rho_{A}\left(\cdot\right).$ We denote by $\rho_{A}$ the density operator of the state $\rho_{A}(\cdot).$
@@
Lemma 1. Let $A,B$ be operators on $H$ satisfying (50), then the overlap of $\rho_{A},\rho_{B}$ is
@@
also $s(A,B)=\frac{1}{4}\left\|A-B\right\|_{2}^{2}.$ By using the estimates (36), (52) we get
@@
For further results concerning the trace-norm distance of fermionic Gaussian states see the recent paper [12].

### v5

@@ 1 Introduction and results
$\left\|\rho_{m_{1},\alpha}-\rho_{{}_{m_{2},\beta}}\right\|_{1}\leq\sqrt{8\min\left(a,b\right)\left\|m\right\|_{2}^{2}+\left\|\alpha-\beta\right\|_{2}^{2}},$ (1)
@@
$2\sqrt{4\min\left(a,b\right)\left\|m_{1}-m_{2}\right\|_{2}^{2}+\frac{1}{2}\left\|\alpha-\beta\right\|_{2}^{2}+\min\left\{P(a,b),P(b,a)\right\}\left\|\alpha-\beta\right\|_{1}},$
@@ A Gaussian state $\rho_{m,\alpha}$ is determined by its quantum characteristic function
$\mathrm{Tr}\sqrt{\rho_{\alpha}}\sqrt{\rho_{m,\beta}}=\frac{\left(\det\hat{\alpha}\det\hat{\beta}\right)^{1/4}}{\det\sigma(\hat{\alpha},\hat{\beta})^{1/2}}\exp\left(-\frac{1}{4}m^{t}\sigma(\hat{\alpha},\hat{\beta})^{-1}m\right),$ (29)
@@
$\exp\left[-\frac{1}{4}\mathrm{Tr\,}d(\alpha,\beta)-\frac{1}{2}m^{t}\left(\hat{\alpha}+\hat{\beta}\right)^{-1}m\right]\leq\mathrm{Tr}\sqrt{\rho_{\alpha}}\sqrt{\rho_{m,\beta}}$ (30)
$\leq$ $\exp\left[-\frac{1}{4}\frac{\ln(1+\left\|d(\alpha,\beta)\right\|)}{\left\|d(\alpha,\beta)\right\|}\mathrm{Tr\,}d(\alpha,\beta)-\frac{1}{2}m^{t}\left(\hat{\alpha}+\hat{\beta}\right)^{-1}m\right].$ (31)
@@
$1-\left(\mathrm{Tr}\sqrt{\rho_{\alpha}}\sqrt{\rho_{m,\beta}}\right)^{2}\leq m^{t}\left(\hat{\alpha}+\hat{\beta}\right)^{-1}m+\frac{1}{2}\mathrm{Tr\,}d(\alpha,\beta).$ (32)
@@
$\left\|\rho_{\alpha}-\rho_{{}_{m,\beta}}\right\|_{1}\leq 2\sqrt{m^{t}\left(\hat{\alpha}+\hat{\beta}\right)^{-1}m+\frac{1}{8}\mathrm{Tr}\left(\hat{\alpha}-\hat{\beta}\right)\left(\hat{\beta}^{-1}-\hat{\alpha}^{-1}\right)}.$ (33)
@@ 3.1 Pure states
$\left\|\rho_{\alpha}-\rho_{{}_{m,\beta}}\right\|_{1}\leq 2\sqrt{\frac{1}{2}m^{t}\left(\alpha+\beta\right)^{-1}m+\frac{1}{4}\mathrm{Tr}\left(\alpha-\beta\right)^{2}}.$
@@
$\left\|\rho_{\alpha}-\rho_{{}_{m,\beta}}\right\|_{1}\leq\sqrt{8\min\left(a,b\right)\left\|m\right\|_{2}^{2}+\left\|\alpha-\beta\right\|_{2}^{2}},$ (39)
@@ 3.2 Gauge-invariant states
$\frac{1}{2}\mathrm{Tr\,}d(\alpha,\beta)=\frac{1}{8}\mathrm{Tr}\left(\hat{\alpha}-\hat{\beta}\right)\left(\hat{\beta}^{-1}-\hat{\alpha}^{-1}\right)$
$\leq\frac{1}{2}\left\|\alpha-\beta\right\|_{2}^{2}+\frac{1}{2}\left\|\alpha\Upsilon_{\alpha}-\beta\Upsilon_{\beta}\right\|_{2}^{2}.$ (45)
@@
$=$ $P(a,b)\left\|\beta-\alpha\right\|_{1}.$
By using (45), (32) and (38) with $\alpha,\beta$ replaced by $\hat{\alpha},\hat{\beta}$, we obtain the estimate (3) for the trace-norm distance of two Gaussian states.
@@ 5 Appendix: the case of fermionic Gaussian states of CAR
and it is called covariance operator of the the state $\rho\left(\cdot\right).$Conversely, any such operator defines a state according to (50), (49) a quasifree state which is denoted $\rho_{A}\left(\cdot\right).$ We denote by $\rho_{A}$ the density operator of the state $\rho_{A}(\cdot).$
@@
Lemma 1. Let $A,B$ be operators in $H$ satisfying (50), then the overlap of $\rho_{A},\rho_{B}$ is
@@
also $s(A,B)=\frac{1}{4}\left\|A-B\right\|_{2}^{2}.$ By using the estimate (37) we get
@@
For further results concerning the trace-norm distance of fermionic Gaussian states cf. recent paper [12].